OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation y=2 when x=3

OpenStudy (anonymous):

simple.. start with y=a/x.. next substitute values and find what a is..!

OpenStudy (anonymous):

y=x/6 ?

OpenStudy (akshay_budhkar):

wrong lyssah.. its inversely propotional what you have is directly propotional...

OpenStudy (anonymous):

y=6/x ?

OpenStudy (akshay_budhkar):

right...

OpenStudy (anonymous):

you just put the value of a in that equation !!